OpenStudy (anonymous):

Please solve x2-x-30/x-5=11 and please show work.

OpenStudy (anonymous):

answer is \[6\pm\sqrt{11}\]

OpenStudy (anonymous):

here is the work

OpenStudy (anonymous):

\[\frac{x^2-x-30}{x-5}=11\] \[x^2-x-30=11(x-5)\] \[x^2-x-30=11x-11\] \[x^2-12x-19=0\]

OpenStudy (anonymous):

now you have a choice, quadratic formula or completing the square. will give the same answer but since 12 is even i choose to complete the square

OpenStudy (anonymous):

typo \[x^2-x-30=11x-55\] \[x^2-12x+15=0\]

OpenStudy (anonymous):

still i complete the square [x^2-12x=-15\] \[(x-6)^2=-15+6^2\] \[(x-6)^2=11\] \[x-6=\pm\sqrt{11}\] \[x=6\pm\sqrt{11}\]

OpenStudy (anonymous):

for 11(x-5 ) how did you get 11x-11 if 11x5 is 55 ?

OpenStudy (anonymous):

that was a mistake that i corrected further down. sorry. where i wrote "typo" i corrected that mistake

OpenStudy (anonymous):

oh i see it. thanks alot :)

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

x2-x-30/x-5=11 x2-x-30 = 11(-x+5) x2-x-30 = -11x+ 55 x2+10x=55+30 x2+10x=85 x2+10x-85=0 \[\Delta = B ^{2}-4ac\] 100x-4(1*85) 100x-340 x= -340/100 x= -3,4

OpenStudy (anonymous):

fraid that is not right

OpenStudy (anonymous):

how did you get 85?

OpenStudy (anonymous):

Im suppose to get one answer. how do i know which either one is right?